OpenStudy (anonymous):

The number of traffic violations in a city increased from 12 to 16 in a month. What is the percent of increase in the number of traffic violations for that month? Round to the nearest percent, if necessary

OpenStudy (gorv):

inc amount /total new amount=(16-12)/16=4/16=1/4 for %=1/4 *100=25%

OpenStudy (anonymous):

% change = (100%)(new-old)/old = (100%)(16-12)/12 = 4/12 = 33%

OpenStudy (anonymous):

The population in a small city increased from 1250 people to 2780 people. What is the percent of increase in the number of people? Round to the nearest percent.